Эz - С 2xth Зҳtsy = |
Answer #1


2= 3 3x + 5y = 1 1-sy put this ain 2x +49 - 22=0 2(t-sy +49 - 22 = 0 0 z e o 2-loy + 129 - 6 2 = 2 +2y =62 2 +2y =62 2 try -